OpenStudy (britbrat1997):

The sum of three consecutive natural numbers is 606, find the numbers. The three numbers in increasing order are ___,___ and ___

OpenStudy (solomonzelman):

The sum of three consecutive integers is 18, find the numbers. Just in case you didn't know, \(«\)integers\(»\) are numbers from \(1\) and on ... The few first integers are: \(1\), \(2\), \(3\), \(4\), \(5\), \(6\), \(7\), \(...\) So, you can let your first integer to be \(x\), and you can write the sum of three consecutive integers as: \(\color{#000000}{ \displaystyle x+x+1+x+2 }\) and sicne you know that the sum of three consecutive integers i equal to 18, you can write the following equation. \(\color{#000000}{ \displaystyle x+x+1+x+2=18 }\) If you add like terms, you get, \(\color{#000000}{ \displaystyle 3x+3=18 }\) Then, you can solve this for \(x\). (Rememeber that \(x\) is your first integer. \(\color{#000000}{ \displaystyle 3x+3=18 }\) \(\color{#000000}{ \displaystyle 3x+3\color{red}{-3}=18\color{red}{-3} }\) \(\color{#000000}{ \displaystyle 3x=15 }\) \(\color{#000000}{ \displaystyle 3x\color{red}{\div3}=15\color{red}{\div3} }\) \(\color{#000000}{ \displaystyle x=5 }\) So, your first integer is \(5\), and thus the other following two are: \(6\) and \(7\). NOTICE, that indeed \(5+6+7=18\).